A new smartphone can be purchased for $840. The phone’s value has a half-life of 23 months and can be modelled by the function 

𝑉 = 840(0.5) 𝑡/23

Algebraically determine how long, to the nearest tenth of a month, it takes the smartphone to be worth $601



1
Expert's answer
2021-11-25T09:14:11-0500

"V=840(0.5)^\frac {t} {23}"


Given "V=601"


"601=840(0.5)^\frac {t} {23}"


"\frac {601}{840}=(0.5)^\frac {t} {23}"


"Log(\frac {601} {840}) =\frac {t} {23} Log 0.5"


"-0.145404814=\frac {t} {23} ×-0.301029996"


"t=\frac {23×0.145404814}{0.301029996}"


"t=11.1095"


"t=11.11months"